# SWAT-NN: Simultaneous Weights and Architecture Training for Neural Networks in a Latent Space

### v2

@@ Abstract
Designing neural networks typically relies on manual trial and error or a neural architecture search (NAS) followed by weight training. The former is time-consuming and labor-intensive, while the latter often discretizes architecture search and weight optimization. In this paper, we propose a fundamentally different approach that simultaneously optimizes both the architecture and the weights of a neural network. Our framework first trains a universal multi-scale autoencoder that embeds both architectural and parametric information into a continuous latent space, where functionally similar neural networks are mapped closer together. Given a dataset, we then randomly initialize a point in the embedding space and update it via gradient descent to obtain the optimal neural network, jointly optimizing its structure and weights. The optimization process incorporates sparsity and compactness penalties to promote efficient models. Experiments on synthetic regression tasks demonstrate that our method effectively discovers sparse and compact neural networks with strong performance.
Index Terms:
Neural Network Optimization, Functional Embedding, Autoencoder, Sparse Networks, Neural Architecture Search, Latent Space Optimization
@@ I Introduction
To address this challenge, Neural Architecture Search (NAS) has emerged as a promising framework for automatically discovering effective neural network architectures [1, 2]. Existing NAS methods can be broadly classified into discrete and continuous search approaches. Early NAS methods predominantly employed Reinforcement Learning(RL) [3] and evolutionary algorithms [4, 5], which fall under the discrete search category. RL-based methods [6, 7] formulate architecture design as a sequential decision-making process, where each component of the architecture is treated as an action, and the objective is to find a sequence of actions that maximizes performance. Alternatively, evolutionary algorithms [8, 9, 10] search by iteratively mutating and recombining candidate operations to evolve better-performing architectures. Despite their differences, both approaches operate over high-dimensional discrete representations and navigate a vast combinatorial search space defined by a predefined set of operations. In contrast to discrete strategies, continuous NAS methods [2, 11] have gained attention for their efficiency and ability to operate in a differentiable search space[12, 13, 14]. These methods create a continuous (e.g., softmax) version of the decision objective, enabling search via gradient descent. The architecture is typically modeled as a computational graph, where each edge carries a soft weight over a predefined set of candidate operations. Nevertheless, while continuous NAS methods significantly reduce search cost and enable differentiable optimization, they are often constrained by task-specific predictors or datasets, and typically search for architectures and weights in a decoupled or alternating manner.
@@
In this work, we propose SWAT-NN, which jointly optimizes both the architecture and its corresponding weights to achieve high performance on given datasets. It is based on a multi-scale autoencoder framework, trained to embed functionally similar neural networks close to one another within a continuous latent space. The encoder learns to represent the full neural network – including both its structure and parameters – in an embedding space, from which multiple decoders generate functionally similar neural networks of varying depths. This embedding, once trained, enables gradient-based optimization (with respect to the embedding, instead of weights) by directly optimizing task-specific performance. Because the latent representation jointly encodes architecture and weights, our method performs unified optimization over both, without requiring separate architectural parameters or surrogate models. We evaluate our proposed approach in the context of multi-layer perceptrons (MLPs) with three different activation functions: sigmoid, tanh, and leaky ReLU, and apply it to the Continuous Optimization Benchmark Suite from Neural Network Regression (CORNN) dataset, which contains 54 regression tasks derived from diverse benchmark functions [15]. Notably, the proposed approach diverges from conventional NAS paradigms by treating networks holistically as function approximators, rather than representing them as graphs composed of discrete operations. This allows the embedding to be over their actual functional behavior. Our main contributions can be summarized as follows:
@@
    Through extensive experiments on 54 regression tasks, we show that SWAT-NN discovers significantly sparser and more compact models compared to existing methods, while maintaining comparable or better accuracy.
@@ II Related Work
Neural networks can be designed manually or through automated search techniques known as Neural Architecture Search (NAS). Traditional NAS approaches include reinforcement learning (RL) [3] or evolutionary algorithms [4, 5]: RL formulates architecture design as a sequential decision-making process, whereas evolutionary methods construct architectures by iteratively mutating or recombining predefined operations. However, these methods typically operate in a discrete search space, making them computationally expensive and difficult to scale. More recent efforts explore continuous NAS, which relaxes the discrete architecture space into a continuous domain.
@@
Differentiable Architecture Search (DARTS) [11] is one of the most influential gradient-based NAS methods. It relaxes discrete architectural choices by introducing a softmax over all candidate operations on each edge of a directed acyclic graph (DAG). The search process is framed as a bi-level optimization problem, in which the network weights are optimized on the training set, while the architecture parameters are updated on the validation set. Such decoupling breaks joint optimization and can result in suboptimal architectures due to misaligned objectives between weight and architecture updates.
@@
Neural Architecture Optimization (NAO) based on a graph variational autoencoder (VAE) [16] encodes network architectures into a continuous latent representation specific to the given set of datasets. It uses a surrogate performance predictor, a regression model estimating task-specific accuracy of an architecture based on its latent representation, to guide the search. Specifically, it performs gradient ascent in the latent space to find architectures that are predicted to perform well for the given dataset. While NAO improves the smoothness of the optimization landscape and captures mathematical properties of architectures through its encoder, it fully decouples architecture discovery from weight optimization and relies on dataset-specific autoencoder and predictors.
@@
In contrast, we propose a novel approach that departs fundamentally from conventional NAS formulations. Rather than searching over discrete or graph-based compositions of operations, we treat entire neural networks as function approximators and embed their complete mathematical behavior – including both architecture and weights – into a universal continuous latent space. Optimization is performed directly on this functional representation, without relying on manually defined search spaces or architectural priors. Moreover, our method is agnostic to specific datasets, as it does not depend on task-specific performance predictors or surrogate models, enabling more general and flexible network discovery. Unlike many the existing methods, which generate the networks through a sequence of decisions [6, 7, 16, 4, 5], our approach decodes a complete neural network from a point in the embedding space.
@@ III-A Embedding MLPs with Autoencoder
We aim to encode both the architecture and corresponding weights of a neural network into a shared embedding space [17]. This is achieved through a multi-scale autoencoder framework. The encoder, denoted as $E$, maps a given MLP into a latent representation. A set of $L$ decoders, denoted as ${D_{1},D_{2},\dots,D_{L}}$, then, generate MLPs with 1 to $L$ hidden layers, respectively. The autoencoder is trained to generate MLPs that may not be reconstructions of the input networks, but are functionally similar – that is, they produce similar outputs when given the same inputs. In the following, we provide further details on the MLP representation and autoencoder design:
@@ III-A2 Multi-scale Autoencoder
Figure 2: Left: Multi-scale autoencoder architecture with four decoders, each corresponding to MLPs with 1–4 hidden layers. Right: Pipeline for training optimal MLPs through gradient-based optimization in the continuous latent space learned by the autoencoder.
The encoder and decoders are implemented using a GPT-2 architecture [18]. Following the matrix representation of MLPs described in Section III-A1, each row $h$ corresponds to the concatenation of all outgoing weights from the $h$th neuron across every layer of the MLP. In other words, the $h$th row aggregates the weights from the $h$th neuron in each layer to all neurons in the subsequent layers, providing a unified view of its outgoing connections throughout the network. Accordingly, we treat the set of outgoing weights from the $h$th neuron across all layers as a single unified token. Each row $h$ is, then, linearly projected to be compatible with the token embedding size of the GPT-2 encoder and decoders. This representation bounds the number of tokens by the maximum number of neurons per layer, thereby minimizing computational cost. By concatenating outgoing weights across layers, the model is further enabled to capture long-range dependencies.
@@ III-A3 Varying Number of Neurons
To enable the autoencoder to account for different numbers of neurons per layer as part of the network architecture, we introduce a neuron indicator matrix to explicitly encode this structural variation. For an MLP with $L$ hidden layers, each with at most $N$ neurons, we augment the matrix representation introduced in Section III-A1 by appending an additional mask matrix $M\in\{0,1\}^{N\times L}$. Each column of $M$ corresponds to one hidden layer, which indicates the activation status of all the neurons in that layer. Specifically, $M_{h,j}=1$ if the $h$th neuron in the $j$th hidden layer is active in the input MLP, and $M_{h,j}=0$ otherwise.
@@ III-A5 Training Autoencoder
We generate input vectors $x$ , and feed them into the input MLP $\mathcal{N}_{s}$ to obtain the corresponding outputs. Each input MLP is then passed through the autoencoder to produce decoded MLPs with varying depths $\{1,2,\dots,L\}$. We adopt a minimum-loss objective to encourage functional similarity between the input MLP and its decoded variants [17]. The loss function is defined as:
@@
where $\mathcal{N}_{s}(x)$ denotes the output of the original input MLP evaluated at input $x$, and $[D_{i}(E(\mathcal{N}_{s}))](x)$ denotes the output of the decoded MLP with $i$ hidden layers, reconstructed from the embedding and evaluated at the same input $x$. This objective encourages at least one decoded MLP to closely approximate the functional behavior of the original MLP across all sampled inputs. It effectively pulls functionally similar MLPs closer together in the embedding space.
The runtime complexity for training a single encoder and $L$ decoders is $\mathcal{O}\left(E\cdot\frac{N}{B}\cdot L\cdot f\right)$, where $f$ is the cost of a forward and backward pass through GPT-2, $N$ is the dataset size, $B$ is the batch size, and $E$ is the number of training epochs.
@@ III-B1 Training Optimal MLPs for a Dataset
where $[D_{i}(z)](x)$ denotes the output of the decoded MLP with $i$ hidden layers evaluated at input data $x$. The gradient of $L_{i}$ with respect to $z$ can be written as:
@@
where $\eta$ is the learning rate and $t\in\{0,1,\dots,T\}$ denotes the optimization step. The runtime complexity for this optimizing is $\mathcal{O}\left(L\cdot T\cdot f\right)$, where $L$ is the number of decoders, $T$ is the number of gradient steps per decoder, and $f$ is the cost of a forward and backward pass through a GPT-2 decoder.
@@ III-B2 Sparsity and Compactness
Here, $M_{i}$ denotes the soft neuron activation indicators for the $i$th hidden layer, and $L$ is the number of hidden layers. The first term encourages the activation probabilities within each layer to become more polarized, while the second term penalizes the average activation level to further suppress unnecessary neurons. Coefficients $\alpha$ and $\beta$ are hyperparameters controlling the strength of each term. During training, neuron indicators smaller than a fixed threshold $t_{n}=0.5$ are softly masked using a sharpened sigmoid function that closely approximates hard thresholding while maintaining differentiability; at test time, neurons this threshold are hard-masked as inactive.
@@ III-B3 Activation Function Selection
We adopt a temperature-controlled softmax mechanism [11] to gradually fix the activation function assigned to each neuron. We define a temperature schedule over training epochs as follows:
@@
Here, $T_{\text{init}}$ is the temperature at the beginning of training, $T_{\text{final}}$ is the minimum temperature reached. $E_{\text{anneal}}$ is the number of epochs over which the temperature linearly decays to $T_{\text{final}}$, and $e$ denotes the current training epoch.
During training, for each neuron, a softmax over activation function candidates (i.e., Leaky ReLU, Tanh, Sigmoid) is applied using the current temperature $T(e)$. Let $\mathbf{a}_{h,j}\in\mathbb{R}^{A}$ denote the unnormalized logits for $A$ activation candidates of the $h$th neuron in layer $j$; the contribution weight of activation operation $a$ at neuron $(h,j)$ is computed as:
$\alpha_{\text{(h,j),k}}=\frac{\exp\left(\mathbf{a}_{(h,j),k}/T(e)\right)}{\sum\limits_{k=1}^{A}\exp\left(\mathbf{a}_{(h,j),k}/T(e)\right)}$ (15)
@@ IV Experiments and Results
  • Exp 1: Learning Sparse and Compact MLPs (Joint Optimization of Activation, Neuron Usage, and Sparsity). We evaluate SWAT-NN, which simultaneously optimizes activation functions, neuron usage, and weight sparsity. We compare against DARTS, a baseline that supports differentiable architecture search, followed by Alternating Direction Method of Multipliers (ADMM) pruning [19].
@@
All the code is publicly available  [^1] https://tinyurl.com/SWAT-NN.
@@ IV-1 Dataset
All evaluations are conducted on the CORNN dataset [15], which consists of 54 synthetic regression tasks. Each task defines a distinct function $f:\mathbb{R}^{2}\rightarrow\mathbb{R}$, designed to capture a wide range of functional behaviors, including periodicity, discontinuities, sharp curvature, and varying smoothness. For each task, the data is split into 3,750 training samples and 1,250 test samples. All inputs are normalized to lie within $[-1,1]^{2}$, and outputs are scaled to the range $(-1,1)$.

### v3

@@ Abstract
Designing neural networks typically relies on manual trial and error or a neural architecture search (NAS) followed by weight training. The former is time-consuming and labor-intensive, while the latter often discretizes architecture search and weight optimization. In this paper, we propose a fundamentally different approach that simultaneously optimizes both the architecture and the weights of a neural network. Our framework first trains a universal multi-scale autoencoder that embeds both architectural and parametric information into a continuous latent space, where functionally similar neural networks are mapped closer together. Given a dataset, we then randomly initialize a point in the embedding space and update it via gradient descent to obtain the optimal neural network, jointly optimizing its structure and weights. The optimization process incorporates sparsity and compactness penalties to promote efficient models. Experiments on regression datasets demonstrate that our method effectively discovers sparse and compact neural networks with strong performance.
@@ I Introduction
To address this challenge, Neural Architecture Search (NAS) has emerged as a promising framework for automatically discovering effective neural network architectures [ren2021comprehensive, elsken2019neural]. Existing NAS methods can be broadly classified into discrete and continuous search approaches. Early NAS methods predominantly employed Reinforcement Learning(RL) [zoph2016neural] and evolutionary algorithms [liu2021survey, stanley2002evolving], which fall under the discrete search category. RL-based methods [baker2016designing, zhong2018practical] formulate architecture design as a sequential decision-making process, where each component of the architecture is treated as an action, and the objective is to find a sequence of actions that maximizes performance. Alternatively, evolutionary algorithms [stanley2009hypercube, yao1999evolving, chen2019renas] search by iteratively mutating and recombining candidate operations to evolve better-performing architectures. Despite their differences, both approaches operate over high-dimensional discrete representations and navigate a vast combinatorial search space defined by a predefined set of operations. In contrast to discrete strategies, continuous NAS methods [elsken2019neural, liu2018darts] have gained attention for their efficiency and ability to operate in a differentiable search space[santra2021gradient, xie2018snas, cai2018proxylessnas]. These methods create a continuous (e.g., softmax) version of the decision objective, enabling search via gradient descent. The architecture is typically modeled as a computational graph, where each edge carries a soft weight over a predefined set of candidate operations. Nevertheless, while continuous NAS methods significantly reduce search cost and enable differentiable optimization, they are often constrained by task-specific predictors or datasets, and typically search for architectures and weights in a decoupled or alternating manner.
@@
In this work, we propose SWAT-NN, which jointly optimizes both the architecture and its corresponding weights to achieve high performance on given datasets. It is based on a multi-scale autoencoder framework, trained to embed functionally similar neural networks close to one another within a continuous latent space. The encoder learns to represent the full neural network – including both its structure and parameters – in an embedding space, from which multiple decoders generate functionally similar neural networks of varying depths. This embedding, once trained, enables gradient-based optimization (with respect to the embedding, instead of weights) by directly optimizing task-specific performance. Because the latent representation jointly encodes architecture and weights, our method performs unified optimization over both, without requiring separate architectural parameters or surrogate models. We evaluate our proposed approach in the context of multi-layer perceptrons (MLPs) with three different activation functions: sigmoid, tanh, and leaky ReLU, and apply it to the Continuous Optimization Benchmark Suite from Neural Network Regression (CORNN) benchmark, which contains 54 regression datasets derived from diverse benchmark function datasets [malan2022continuous]. Notably, the proposed approach diverges from conventional NAS paradigms by treating networks holistically as function approximators, rather than representing them as graphs composed of discrete operations. This allows the embedding to be over their actual functional behavior. Our main contributions can be summarized as follows:
@@
    Through experiments on 54 regression datasets, we show that SWAT-NN discovers significantly sparser and more compact models compared to existing methods, while maintaining comparable or better accuracy.
@@ II Related Work
Neural networks can be designed manually or through automated search techniques known as Neural Architecture Search (NAS). Traditional NAS approaches include reinforcement learning (RL) [zoph2016neural] or evolutionary algorithms [liu2021survey, stanley2002evolving]: RL formulates architecture design as a sequential decision-making process, whereas evolutionary methods construct architectures by iteratively mutating or recombining predefined operations. However, these methods typically operate in a discrete search space, making them computationally expensive and difficult to scale. More recent efforts explore continuous NAS, which relaxes the discrete architecture space into a continuous domain.
@@
Differentiable Architecture Search (DARTS) [liu2018darts] is one of the most influential gradient-based NAS methods. It relaxes discrete architectural choices by introducing a softmax over all candidate operations on each edge of a directed acyclic graph (DAG). The search process is framed as a bi-level optimization problem, in which the network weights are optimized on the training set, while the architecture parameters are updated on the validation set. Such decoupling breaks joint optimization and can result in suboptimal architectures due to misaligned objectives between weight and architecture updates.
@@
Neural Architecture Optimization (NAO) based on a graph variational autoencoder (VAE) [li2020ngae] encodes network architectures into a continuous latent representation specific to the given set of datasets. It uses a surrogate performance predictor, a regression model estimating task-specific accuracy of an architecture based on its latent representation, to guide the search. Specifically, it performs gradient ascent in the latent space to find architectures that are predicted to perform well for the given dataset. While NAO improves the smoothness of the optimization landscape and captures mathematical properties of architectures through its encoder, it fully decouples architecture discovery from weight optimization and relies on dataset-specific predictors. Further, it requires retraining of the autoencoder from scratch for each new set of tasks.
@@
In contrast, we propose a novel approach that departs fundamentally from conventional NAS formulations. Rather than searching over discrete or graph-based compositions of operations, we treat entire neural networks as function approximators and embed their complete mathematical behavior – including both architecture and weights – into a universal continuous latent space. Optimization is performed directly on this functional representation, without relying on manually defined search spaces or architectural priors. Moreover, our method is agnostic to specific datasets, as it does not depend on task-specific performance predictors or surrogate models, enabling more general and flexible network discovery. Unlike many the existing methods, which generate the networks through a sequence of decisions [baker2016designing, zhong2018practical, li2020ngae, liu2021survey, stanley2002evolving], our approach decodes a complete neural network from a point in the embedding space.
@@ III-A Embedding MLPs with Autoencoder
We aim to encode both the architecture and corresponding weights of a neural network into a shared embedding space. This is achieved through a multi-scale autoencoder framework. The encoder, denoted as $E$, maps a given MLP into a latent representation. A set of $L$ decoders, denoted as $D_{1},D_{2},\dots,D_{L}$}, then, generate MLPs with 1 to $L$ hidden layers, respectively. In this work, we set $L=4$. The autoencoder is trained to generate MLPs that may not be reconstructions of the input networks, but are functionally similar – that is, they produce similar outputs when given the same inputs. In the following, we provide further details on the MLP representation and autoencoder design:
@@ III-A2 Multi-scale Autoencoder
Figure 2: Left: Multi-scale autoencoder architecture with four decoders, each corresponding to MLPs with 1–4 hidden layers. Right: Pipeline for training optimal MLPs through gradient-based search in the continuous latent space learned by the autoencoder.
The encoder and decoders are implemented using a GPT-2 architecture [radford2019language]. Following the matrix representation of MLPs described in Section III-A1, each row $h$ corresponds to the concatenation of all outgoing weights from the $h$th neuron across every layer of the MLP. In other words, the $h$th row aggregates the weights from the $h$th neuron in each layer to all neurons in the subsequent layers, providing a unified view of its outgoing connections throughout the network. Accordingly, we treat the set of outgoing weights from the $h$th neuron across all layers as a single unified token. Each row $h$ is, then, linearly projected to be compatible with the token embedding size of the GPT-2 encoder and decoders. This representation bounds the number of tokens by the maximum number of neurons per layer, thereby minimizing computational cost. By concatenating outgoing weights across layers, the model is further enabled to capture long-range dependencies.
@@ III-A3 Varying Number of Neurons
To enable the autoencoder to account for different numbers of neurons per layer as part of the network architecture, we introduce a neuron indicator matrix to explicitly encode this structural variation. For an MLP with $L$ hidden layers, each with at most $N$ neurons, we augment the matrix representation introduced in Section III-A2 by appending an additional mask matrix $M\in\{0,1\}^{N\times L}$. Each column of $M$ corresponds to one hidden layer, which indicates the activation status of all the neurons in that layer. Specifically, $M_{h,j}=1$ if the $h$th neuron in the $j$th hidden layer is active in the input MLP, and $M_{h,j}=0$ otherwise.
@@ III-A5 Training Autoencoder
We generate input vectors $x$ , and feed them into the input MLP $\mathcal{N}_{s}$ to obtain the corresponding outputs. Each input MLP is then passed through the autoencoder to produce decoded MLPs with varying depths $\{1,2,\dots,L\}$. To capture the diverse functional properties of neural networks, we train the autoencoder on a large-scale collection of synthetically generated MLPs spanning a wide range of architectures and behaviors. We adopt a minimum-loss objective to encourage functional similarity between the input MLP and its decoded variants. The loss function is defined as:
@@
where $\mathcal{N}_{s}(x)$ denotes the output of the original input MLP evaluated at input $x$, and $D_{i}(E(\mathcal{N}_{s}))](x)$ denotes the output of the decoded MLP with $i$ hidden layers, reconstructed from the bottleneck embedding and evaluated at the same input $x$. This objective encourages at least one decoded MLP to closely approximate the functional behavior of the original MLP across all sampled inputs. It effectively pulls functionally similar MLPs closer together in the embedding space.
The runtime complexity for training a single encoder and $L$ decoders is $\mathcal{O}\left(E\cdot\frac{N}{B}\cdot L\cdot f\right)$, where $f$ is the cost of a forward or backward pass through GPT-2, $N$ is the dataset size, $B$ is the batch size, and $E$ is the number of training epochs.
@@ III-B1 Training Optimal MLPs for a Dataset
where $D_{i}(z)](x)$ denotes the output of the decoded MLP with $i$ hidden layers evaluated at input data $x$. The gradient of $L_{i}$ with respect to $z$ can be written as:
@@
where $\eta$ is the learning rate and $t\in\{0,1,\dots,T\}$ denotes the optimization step. The runtime complexity for this optimizing is $\mathcal{O}\left(L\cdot T\cdot f\right)$, where $L$ is the number of decoders, $T$ is the number of gradient steps per decoder, and $f$ is the cost of a forward or backward pass through a GPT-2 decoder.
@@ III-B2 Sparsity and Compactness
Here, $M_{i}$ denotes the soft neuron activation indicators for the $i$th hidden layer, and $L$ is the number of layers. The first term encourages the activation probabilities within each layer to become more polarized, while the second term penalizes the average activation level to further suppress unnecessary neurons. Coefficients $\alpha$ and $\beta$ are hyperparameters controlling the strength of each term. During training, neuron indicators smaller than a fixed threshold $t_{n}=0.5$ are softly masked using a sharpened sigmoid function that closely approximates hard thresholding while maintaining differentiability; at test time, neurons this threshold are hard-masked as inactive.
@@ III-B3 Activation Function Selection
We adopt a temperature-controlled softmax mechanism [liu2018darts] to gradually fix the activation function assigned to each neuron. We define a temperature schedule over training epochs as follows:
@@
Here, $T_{\text{init}}$ is the initial temperature at the beginning of training, $T_{\text{final}}$ is the minimum temperature reached. $E_{\text{anneal}}$ is the number of epochs over which the temperature linearly decays to $T_{\text{final}}$, and $e$ denotes the current training epoch.
During training, for each neuron, a softmax over activation function candidates (i.e., ReLU, Tanh, Sigmoid) is applied using the current temperature $T(e)$. Let $\mathbf{a}_{h,j}\in\mathbb{R}^{A}$ denote the unnormalized logits for $A$ activation candidates of the $h$th neuron in layer $j$; the contribution weight of activation operation $a$ at neuron $(h,j)$ is computed as:
$\alpha_{\text{h,j,a}}=\frac{\exp\left(\mathbf{a}_{h,j,a}/T(e)\right)}{\sum\limits_{a=1}^{A}\exp\left(\mathbf{a}_{h,j,a}/T(e)\right)}$ (15)
@@ IV Experiments and Results
  • Exp 1: Learning Sparse and Compact MLPs (Joint Optimization of Activation, Neuron Usage, and Sparsity). We evaluate SWAT-NN’s ability to jointly optimizes activation functions, neuron usage, and weight sparsity. We compare against DARTS, a baseline that supports differentiable architecture search, followed by Alternating Direction Method of Multipliers (ADMM) pruning.
@@
All the code is publicly available  [^1] https://github.com/zitonghuangcynthia/SWAT-NN_code.
@@ IV-1 Dataset
All evaluations are conducted on the CORNN benchmark suite [malan2022continuous], which consists of 54 regression datasets. Each dataset defines a distinct function $f:\mathbb{R}^{2}\rightarrow\mathbb{R}$, designed to capture a wide range of functional behaviors, including periodicity, discontinuities, sharp curvature, and varying smoothness. For each dataset, the data is split into 3,750 training samples and 1,250 test samples. All inputs are normalized to lie within $1,1]^{2}$, and outputs are scaled to the range (1,1)(-1,1).